# Reward Certification for Policy Smoothed Reinforcement Learning

### 11 Dec 2023 v1

@@ Abstract
Reinforcement Learning (RL) has achieved remarkable success in safety-critical areas, but it can be weakened by adversarial attacks. Recent studies have introduced ”smoothed policies” in order to enhance its robustness. Yet, it is still challenging to establish a provable guarantee to certify the bound of its total reward. Prior methods relied primarily on computing bounds using Lipschitz continuity or calculating the probability of cumulative reward above specific thresholds. However, these techniques are only suited for continuous perturbations on the RL agent’s observations and are restricted to perturbations bounded by the $l_{2}$-norm. To address these limitations, this paper proposes a general black-box certification method capable of directly certifying the cumulative reward of the smoothed policy under various $l_{p}$-norm bounded perturbations. Furthermore, we extend our methodology to certify perturbations on action spaces. Our approach leverages $f$-divergence to measure the distinction between the original distribution and the perturbed distribution, subsequently determining the certification bound by solving a convex optimisation problem. We provide a comprehensive theoretical analysis and run sufficient experiments in multiple environments. Our results show that our method not only improves the certified lower bound of mean cumulative reward but also demonstrates better efficiency than state-of-the-art techniques.
@@ Introduction
The utilisation of neural networks in Reinforcement Learning (RL) has achieved remarkable success in safety-critical domains, such as controlling robots and autonomous driving (Sallab et al. 2017; Pan et al. 2017; Johannink et al. 2019). Nevertheless, recent research has revealed their vulnerability to the presence of adversarial perturbations (Szegedy et al. 2014; Madry et al. 2017; Goodfellow, Shlens, and Szegedy 2015; Mu et al. 2021; Jin et al. 2022; Zhang et al. 2021). For example, numerous studies have demonstrated that even well-trained RL policies can suffer significant failures when directly perturbing the observations of the RL agent (Pattanaik et al. 2018; Mu et al. 2023) or in action space (Lin et al. 2017). In this regard, it is vital to analyse their robustness and reliability before their deployment in various safety-critical systems (Christiano et al. 2016; Cheng et al. 2019).
Various empirical defences have been proposed to defend adversarial attacks in RL system (Manikandan et al. 2011; Pattanaik et al. 2017), while it has been demonstrated that even robust models can still be compromised by more advanced attack methods (Russo and Proutiere 2019). Hence, there is a need for computing provable guarantees for the trained policy to disrupt the repeated game between attacks and defence, which is referred to as robustness certification. The majority of studies aiming to provide robustness certification primarily focus on classification tasks, while the certification for RL is seldom touched.
Compared to classification tasks, certifying RL engages more challenges due to its sequential decision-making nature. To address this obstacle, recent efforts have focused on certifying a ”smoothed policy” based randomised smoothing strategies. What sets these approaches apart is that they do not require access to the internal architecture and parameters of the DNNs, which is referred to as black-box certification. For instance, Wu et al. (2021) employed Lipschitz continuity to approximate the final output, but this approach resulted in a loose bound. Instead of directly certifying the lower bound, (Kumar, Levine, and Feizi 2021) proposed certifying the reward by breaking it down into several thresholds and then assessing the probability of the cumulative reward staying above a specific threshold. However, this method proved too restrictive (see discussion in Section Comparison with prior work), leading to a loss of essential information regarding the output reward and resulting in weaker certificates for smoothed policies. Additionally, both of these approaches can only handle observation perturbations under the $l_{2}$ norm bound. However, in the real world, adversaries can perturb both observations and actions within the bounds of various $l_{p}$-norm perturbation constraints.
@@
Our contributions can be summarised as: (i) We propose a novel methodology to directly certify the cumulative reward of the smoothed policy. This approach uses $f$-divergence to gauge the separation between the original distribution and the perturbed distribution. Subsequently, we calculate the certification bound by solving a convex optimisation problem. (ii) Our method is capable to handle perturbations bounded by both the $l_{0}$ and $l_{1}$ norm. This work is the first of its kind to consider the certification of the $l_{0}$-norm bounded perturbation in the action space. (iii) By comparing our method with the previous approach, we theoretically validate that our method can enhance certified robustness by taking into account the distribution of cumulative rewards during sampling. (iv) Through comprehensive experiments, we show that our method outperforms the state-of-the-art methods, producing tighter bounds to $l_{2}$ perturbations. Furthermore, we conduct experiments in various environments to demonstrate our certification performance for $l_{1}$ perturbations in observation and $l_{0}$ perturbations in the action space.
@@
As a result, maximising the problem stated in Eq. 9 can provide us with the lower bound for the mean cumulative reward. ∎
@@
Two existing studies concerning the certification of cumulative rewards in RL are (Kumar, Levine, and Feizi 2021) and (Wu et al. 2021). In (Kumar, Levine, and Feizi 2021), the authors tackle certification by forming it as a classification problem, determining whether the cumulative reward surpasses a specific threshold. They utilise integrals to calculate the total expected reward through the cumulative distribution function (CDF). Conversely, (Wu et al. 2021) directly certifies the lower boundary of the anticipated reward by exploiting the Lipschitz continuity of the smoothed reward. Nevertheless, in comparison to both our work and the methodology proposed in (Kumar, Levine, and Feizi 2021), this approach demonstrates a relatively weak performance, as indicated by the findings in our experimental results provided in the Appendix I. This divergence in performance could be potentially due to their emphasis on certifying per-step perturbations. Consequently, in this article, we will present a theoretical analysis that elucidates why our method outperforms the approach introduced in (Kumar, Levine, and Feizi 2021).
The problem studied in (Kumar, Levine, and Feizi 2021) may be treated as a special case of our methodology. When considering the case of smoothed Gaussian noise i.i.d sampling from distribution $\mathcal{N}(0,\sigma^{2}I_{D})$, the Hockey-Stick divergence can be analytically calculated, which can be expressed as a function of the $l_{2}$-norm, as established in (Dong et al. 2021) (for a detailed proof, please refer to the Appendix D). The Hockey-Stick divergence, denoted as $D_{HS,\lambda}(q||p)$, is defined by the function $f(x)=max(x-\lambda,0)-max(1-\lambda,0)$.
@@
(CDF-based method in optimisation framework) Suppose the reward range is (a, b) and considering $n$ thresholds, denoted as $a<g_{1}<g_{2}<...<g_{n}<b$. Let $\theta_{i}$ represent the lower bound of probability that the total reward obtained by the smoothed policy is above the threshold $g_{i}$. Let f* be the convex conjugate of hockey-stick divergences, we have
@@
In this section, we first demonstrate the experimental results of our method compared with the baseline to certify the cumulative reward under $l_{2}$-norm bounded perturbation. Then, we show the results of certifying the $l_{1}$-norm perturbation on observations and the $l_{0}$-norm perturbation on action space.

### 12 Dec 2023 v2

# ReCePS: Reward Certification for Policy Smoothed Reinforcement Learning

@@ Abstract
Reinforcement Learning (RL) has achieved remarkable success in safety-critical areas, but it can be weakened by adversarial attacks. Recent studies have introduced ”smoothed policies” to enhance its robustness. Yet, it is still challenging to establish a provable guarantee to certify the bound of its total reward. Prior methods relied primarily on computing bounds using Lipschitz continuity or calculating the probability of cumulative reward above specific thresholds. However, these techniques are only suited for continuous perturbations on the RL agent’s observations and are restricted to perturbations bounded by the $l_{2}$-norm. To address these limitations, this paper proposes a general black-box certification method, called ReCePS, which is capable of directly certifying the cumulative reward of the smoothed policy under various $l_{p}$-norm bounded perturbations. Furthermore, we extend our methodology to certify perturbations on action spaces. Our approach leverages $f$-divergence to measure the distinction between the original distribution and the perturbed distribution, subsequently determining the certification bound by solving a convex optimisation problem. We provide a comprehensive theoretical analysis and run sufficient experiments in multiple environments. Our results show that our method not only improves the certified lower bound of the mean cumulative reward but also demonstrates better efficiency than state-of-the-art methods.
@@ Introduction
The utilisation of neural networks in Reinforcement Learning (RL) has achieved remarkable success in safety-critical domains, such as controlling robots and autonomous driving (Sallab et al. 2017; Pan et al. 2017; Johannink et al. 2019). Nevertheless, recent research has revealed their vulnerability to the presence of adversarial perturbations (Szegedy et al. 2014; Madry et al. 2017; Goodfellow, Shlens, and Szegedy 2015; Mu et al. 2021; Jin et al. 2022; Zhang et al. 2021). For example, numerous studies have demonstrated that even well-trained RL policies can suffer significant failures when directly perturbing the observations of the RL agent (Pattanaik et al. 2018; Mu et al. 2023) or in action space (Lin et al. 2017). In this regard, it is vital to analyse their robustness before their deployment in safety-critical systems (Christiano et al. 2016; Cheng et al. 2019).
Various empirical defences have been proposed to defend adversarial attacks in RL systems (Manikandan et al. 2011; Pattanaik et al. 2017), while it has been demonstrated that even robust models can still be compromised by more advanced attack methods (Russo and Proutiere 2019). Hence, there is a need for computing provable guarantees for the trained policy to disrupt the repeated game between attacks and defence, which is referred to as robustness certification. The majority of studies aiming to provide robustness certification primarily focus on classification tasks, while the certification for RL is seldom touched.
Compared to classification tasks, certifying RL engages more challenges due to its sequential decision-making nature. To address this obstacle, recent efforts have focused on certifying a ”smoothed policy” based randomised smoothing strategies. What sets these approaches apart is that they do not require access to the internal architecture and parameters of the DNNs, which is referred to as black-box certification. For instance, Wu et al. (2021) employed Lipschitz continuity to approximate the final output, but this approach resulted in a loose bound. Instead of directly certifying the lower bound, (Kumar, Levine, and Feizi 2021) proposed certifying the reward by breaking it down into several thresholds and then assessing the probability of the cumulative reward staying above a specific threshold. However, this method proved too restrictive (see discussion in Section Comparison with prior work), leading to a loss of essential information regarding the output reward and resulting in weaker certificates for smoothed policies. Additionally, both approaches can only handle observation perturbations under the $l_{2}$ norm bound. However, in the real world, adversaries can perturb both observations and actions within the bounds of various perturbation constraints of the $l_{p}$ norm.
@@
Our contributions can be summarised as: (i) We propose a novel methodology to directly certify the cumulative reward of the smoothed policy. This approach uses $f$-divergence to gauge the separation between the original distribution and the perturbed distribution. Subsequently, we calculate the certification bound by solving a convex optimisation problem. (ii) Our method is capable of handling perturbations bounded by both the $l_{0}$ and $l_{1}$ norm. This work is the first of its kind to consider the certification of the $l_{0}$-norm bounded perturbation in the action space. (iii) By comparing our method with the previous approach, we theoretically validate that our method can enhance certified robustness by taking into account the distribution of cumulative rewards during sampling. (iv) Through comprehensive experiments, we demonstrate that our method outperforms the state-of-the-art methods, producing tighter bounds for $l_{2}$ perturbations. Furthermore, we conduct intensive experiments in various environments to demonstrate the effectiveness of our certification method for $l_{1}$ perturbations in observation and $l_{0}$ perturbations in the action space.
@@
As a result, maximising the problem stated in Eq. 9 can give the lower bound for the mean cumulative reward. ∎
@@
Two existing studies concerning the certification of cumulative rewards in RL are (Kumar, Levine, and Feizi 2021) and (Wu et al. 2021). In (Kumar, Levine, and Feizi 2021), the authors tackle certification by forming it as a classification problem, determining whether the cumulative reward surpasses a specific threshold. They utilise integrals to calculate the total expected reward through the cumulative distribution function. Conversely, Wu et al. (2021) directly certifies the lower boundary of the anticipated reward by exploiting the Lipschitz continuity of the smoothed reward. Nevertheless, in comparison to both our work and the methodology proposed by Kumar, Levine, and Feizi (2021), this approach demonstrates a relatively weak performance, as indicated by the findings in our experimental results provided in the Appendix I. This divergence in performance could be potentially due to their emphasis on certifying per-step perturbations. Consequently, in this article, we will present a theoretical analysis that elucidates why our method outperforms the approach introduced in (Kumar, Levine, and Feizi 2021).
The problem studied in (Kumar, Levine, and Feizi 2021) may be treated as a special case of our methodology. When considering the case of smoothed Gaussian noise i.i.d sampling from distribution $\mathcal{N}(0,\sigma^{2}I_{D})$, the Hockey-Stick divergence can be analytically calculated, which can be expressed as a function of the $l_{2}$-norm, as established in (Dong et al. 2021) (detailed proof in Appendix D). The Hockey-Stick divergence, denoted as $D_{HS,\lambda}(q||p)$, is defined by the function $f(x)=max(x-\lambda,0)-max(1-\lambda,0)$.
@@
(CDF-based optimisation framework) Suppose the reward range is (a, b) and considering $n$ thresholds, denoted as $a<g_{1}<g_{2}<...<g_{n}<b$. Let $\theta_{i}$ represent the lower bound of probability that the total reward obtained by the smoothed policy is above the threshold $g_{i}$. Let f* be the convex conjugate of hockey-stick divergences, we have
@@
We first demonstrate the experimental results of our method - ReCePS  [^3] Our code is available at https://github.com/TrustAI/ReCePS, compared with the baseline to certify the cumulative reward under $l_{2}$-norm bounded perturbation. Then, we show the results of certifying the $l_{1}$-norm perturbation on observations and the $l_{0}$-norm perturbation on action space.